solve (xz-y)p + (yz - x)q =1-z^2
1
Expert's answer
2020-12-29T15:11:45-0500

P=zx-y

Q=zy-x

R=1-z2

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{zx-y}=\frac{dy}{zy-x}=\frac{dz}{1-z^2}"

Multiplyers: z,1,x

zdx+dy-xdz=0

"zx+y+zx=2zx+y=C_1"

Multiplyers: 1,z,y

dx+zdy+ydz=0

"x+zy+zy=x+2zy=C_2"

The general solution is,

"\phi(c_{1},c_{2})=0\\ \phi\left(2zx+y, 2zy+x\right)=0"